# Proxy-Free GFlowNet

### v1

@@ Abstract
Generative Flow Networks (GFlowNets) are a promising class of generative models designed to sample diverse, high-reward structures by modeling distributions over compositional objects. In many real-world applications, obtaining the reward function for such objects is expensive, time-consuming, or requires human input, making it necessary to train GFlowNets from historical datasets. Most existing methods adopt a model-based approach, learning a proxy model from the dataset to approximate the reward function. However, this strategy inherently ties the quality of the learned policy to the accuracy of the proxy, introducing additional complexity and uncertainty into the training process. To overcome these limitations, we propose Trajectory-Distilled GFlowNet (TD-GFN), a proxy-free training framework that eliminates the need for out-of-dataset reward queries. Our method is motivated by the key observation that different edges in the associated directed acyclic graph (DAG) contribute unequally to effective policy learning. TD-GFN leverages inverse reinforcement learning to estimate edge-level rewards from the offline dataset, which are then used to ingeniously prune the DAG and guide backward trajectory sampling during training. This approach directs the policy toward high-reward regions while reducing the complexity of model fitting. Empirical results across multiple tasks show that TD-GFN trains both efficiently and reliably, significantly outperforming existing baselines in convergence speed and sample quality.
@@ 1 Introduction
Generative Flow Networks (GFlowNets or GFNs) [1, 2] are generative models designed to sample compositional objects in proportion to their positive rewards, and have shown strong empirical success across diverse domains including molecule discovery [1], biological sequence design [3], combinatorial optimization [4], fine-tuning of large language models [5] and diffusion models [6].
Despite the promising potential of GFlowNets, interacting with the environment to obtain reward feedback for policy updates is often impractical or infeasible in real-world applications—such as biological and chemical tasks involving costly experiments or simulations, or tasks requiring human evaluation [7]. As a result, learning GFlowNets from historically collected datasets becomes essential.
Current approaches typically adopt a model-based strategy, wherein a proxy model is trained on the dataset to approximate the reward signal from the environment for updating the policy [1, 3, 8, 9, 10]. However, reliance on such proxy reward models imposes a fundamental limitation on the performance of the learned policies, as their effectiveness is inherently constrained by the accuracy of the reward approximation. Inaccuracies or biases in the proxy model inevitably propagate through the learning pipeline, potentially resulting in suboptimal or misleading policies [11, 12, 13]. This challenge is particularly pronounced in domains such as language modeling [14, 15], recommendation systems [16], and autonomous driving [17], where surrogate reward models often struggle to generalize beyond the distribution of their training data. Furthermore, constructing a high-quality reward model typically demands a large and diverse dataset, along with substantial domain-specific expertise—both of which are resource-intensive [18, 19, 20]. Consequently, dependence on proxy reward models introduces considerable burden, complexity, and uncertainty into the training process, ultimately undermining the reliability and scalability of the overall algorithm.
The limitations of existing training methods naturally give rise to the following question:
Can we train GFlowNets effectively and efficiently without a proxy reward model?
The absence of a proxy reward model implies that out-of-distribution feedback is unavailable during training—posing a significant challenge for GFlowNets, whose objectives inherently require broad coverage of the state space. Consequently, an effective training framework must fully leverage the available historical data, while also generalizing strategically to sufficiently broad unseen regions of the state space, in order to achieve a stronger balance between exploration and exploitation.
Several recent efforts have attempted to address this challenge. An early approach [21] uses offline action probabilities to constrain the learned policy, while COFlowNet [22] penalizes flow on unseen edges to limit policy coverage to dataset-adjacent regions. Both methods aim to improve policy reliability by imposing coarse constraints that merely align policy behavior or coverage more closely with the dataset. However, such constraints impose blunt limitations that hinder generalization beyond the dataset and make performance highly sensitive to data quality and coverage, as they fail to fully exploit the trajectory-level information contained in the dataset. Additionally, [23] explores an “offline” setting with backward sampling from terminal states, but primarily focuses on evaluating sampling strategies rather than proposing a concrete training framework.
To address these challenges, we begin with a key insight: different edges in the environment’s directed acyclic graph (DAG) contribute unequally to effective GFlowNet training. As shown in Figure 1(a), removing the red edge $(n_{1}\rightarrow n_{2})$ blocks access to a terminal node with reward $20$, while removing the green edge $(n_{4}\rightarrow n_{7})$ only affects a low-reward node ($0.1$). This disparity underscores the varying importance of edges in learning the target distribution (Figure 1(b)), which can in turn be leveraged to guide the policy in selectively exploring unseen regions (Figure 1(c)).
@@
Motivated by this insight, we propose a novel proxy-free GFlowNet training framework, Trajectory-Distilled GFlowNet (TD-GFN), which eliminates reliance on proxy models by avoiding out-of-distribution reward queries. Unlike prior methods that primarily differentiate behaviors based on their proximity to the dataset, TD-GFN provides more informative supervision by capturing the varying importance of transitions across the environment DAG, thereby enabling a more effective exploration–exploitation trade-off.
Specifically, TD-GFN introduces a novel edge reward function that quantifies the contribution of each transition in the environment DAG. This function is inferred via inverse reinforcement learning (IRL) from a carefully rebalanced offline dataset, leveraging the theoretical equivalence between GFlowNet training and entropy-regularized reinforcement learning (RL) [24]. The structured supervision provided by the resulting edge rewards enables the policy to generalize strategically beyond observed trajectories, rather than remain confined to dataset-supported regions.
To mitigate the sensitivity to function approximation errors introduced by directly incorporating edge rewards into the training objective—an issue analogous to the limitations observed with proxy reward models—we propose two indirect mechanisms for shaping the GFlowNet policy. First, we prune the environment DAG by removing low-utility transitions as determined by the learned edge rewards. This yields a more compact yet expressive action space that prioritizes structurally informative and high-reward trajectories, while reducing the complexity of model fitting. Second, we introduce a prioritized backward sampling procedure over the pruned DAG, which reinforces the core inductive bias of GFlowNets—allocating attention in proportion to reward—while improving training efficiency by focusing updates on high-value regions of the state space.
Equipped with these novel components, TD-GFN exhibits strong empirical performance across diverse benchmark tasks. Compared to a broad range of baselines, including proxy-based GFlowNet training paradigms, TD-GFN consistently achieves superior performance and faster convergence, underscoring its potential as a more stable and reliable alternative for efficient GFlowNet training.
In summary, we introduce Trajectory-Distilled GFlowNet (TD-GFN), a novel proxy-free training framework that eliminates the need for proxy reward queries on out-of-distribution objects by extracting structural supervision from existing data. The key contributions of TD-GFN are as follows:
  • TD-GFN employs inverse reinforcement learning to infer edge-level rewards from a rebalanced offline dataset, capturing the relative importance of transitions within the environment DAG. These edge rewards provide structured, transition-level supervision that enables the policy to generalize beyond the dataset while remaining grounded in reliable behavior.
  • Leveraging the inferred edge rewards, TD-GFN prunes low-utility transitions from the environment DAG and introduces a prioritized backward sampling procedure guided by both edge and terminal rewards. This approach concentrates learning on informative regions of the state space and enhances training efficiency by reducing the complexity of model fitting.
  • TD-GFN consistently demonstrates strong empirical performance across diverse benchmark tasks, outperforming a wide range of baselines in both convergence speed and final policy quality. These results highlight its effectiveness as a stable and efficient training paradigm within the GFlowNet framework.
@@ 2 Preliminaries
In this section, we describe Generative Flow Networks (GFlowNets or GFNs) [1, 2] and the maximum causal entropy inverse reinforcement learning (IRL) framework [25, 26], which help clarify our proposed method. Additional related works are provided in Appendix B for completeness.
@@ 2.1 GFlowNets
GFlowNets are defined within an environment represented as a directed acyclic graph (DAG), denoted by $G=(V,E)$, where $V$ is the set of nodes and $E$ the set of directed edges. A node $s^{\prime}$ is said to be a child of node $s$ if and only if there exists a directed edge $(s\rightarrow s^{\prime})\in E$; correspondingly, $s$ is the parent of $s^{\prime}$. The graph includes a unique root node $s_{0}\in V$ with no incoming edges and a subset of terminal nodes $\mathcal{X}\subseteq V$ with no outgoing edges.
A positive reward function $R(x)$ is defined over the terminal nodes $x\in\mathcal{X}$. The objective of GFlowNets is to learn a stochastic forward policy $\mathcal{P}_{F}(s^{\prime}|s)$ such that the probability of reaching a terminal node $x\in\mathcal{X}$ is proportional to its associated reward $R(x)$. This is accomplished by generating sequences of transitions, where at each step, the agent selects a child node of the current state. In this work, we focus on deterministic environment dynamics, where each action uniquely determines the next state by traversing the corresponding edge in the DAG, as is standard in most existing GFlowNet studies [1, 3, 4, 5, 6, 9, 10].
For training GFlowNets, [1] introduced a parameterized flow function and proposed the Flow-Matching (FM) objective, whose global minimization ensures that the induced policy satisfies the desired proportionality condition. Recent work [24] establishes a formal equivalence between GFlowNet training and entropy-regularized RL under specific settings: a discount factor $\gamma=1$, an entropy regularization coefficient $\lambda=1$, and a modified reward function defined by a fixed backward policy $\mathcal{P}_{B}(s|s^{\prime})$ that samples parent nodes in the DAG in reverse from a given child node:
@@ 2.2 Maximum causal entropy IRL
Maximum causal entropy inverse reinforcement learning (IRL) [25, 26] seeks to recover a reward function $r(s,s^{\prime})$ that both explains expert behavior under a policy $\pi_{E}$ and promotes maximum entropy in the learned policy. This objective is formalized as the following minimax optimization problem:
[…]
where $\mathcal{H}(\pi)$ denotes the causal entropy of the policy $\pi$, consistent with its definition in Equation 2. In our approach, we adapt this framework to infer an edge-level reward function $R_{E}:E\rightarrow\mathbb{R}$ that reflects the expert policy’s preferences for allocating flow across specific edges in the graph.
@@ 3 Method
In this section, we introduce Trajectory-Distilled GFlowNet (TD-GFN), a proxy-free training framework that eliminates the need to query rewards outside the dataset, thereby avoiding reliance on proxy reward models. Unlike prior methods that constrain the policy to remain close to the dataset [21, 22], TD-GFN achieves a more effective exploration–exploitation trade-off by strategically allocating attention across the environment DAG.
The overall training pipeline is illustrated in Figure 2. First, TD-GFN learns an edge reward function from a rebalanced dataset using inverse reinforcement learning (IRL). This function captures the relative importance of each edge in the environment DAG, providing principled guidance for policy learning. Leveraging these edge-level rewards, TD-GFN then prunes the DAG to remove low-utility transitions and trains the policy via prioritized backward sampling over the resulting subgraph. This design enables strategic generalization to broad, unseen regions of the state space while maintaining alignment with high-reward behaviors. The following subsections detail each component.
@@
A comprehensive summary of the full training procedure is presented in Algorithm 1 (Appendix A). Notably, the contributions of TD-GFN are orthogonal to the specific training strategies employed after trajectory sampling, as discussed in Appendix D.5.
@@ 3.1 Edge rewards extraction via IRL
Motivated by the insight that different edges in the environment DAG contribute unequally to effective policy learning, we aim to quantify this disparity to enable more targeted and generalizable training. Nevertheless, measuring the contribution of individual transitions under the inherently stochastic nature of GFlowNet policies poses a fundamental challenge. To tackle this, we leverage the theoretical equivalence between GFlowNet training and entropy-regularized reinforcement learning (Equation 2) and applying the maximum causal entropy IRL framework [25, 26] (Equation 3) to extract a fine-grained edge-level reward function from offline trajectories. These learned edge rewards act as structured supervision signals that capture the relative importance of transitions, guiding the policy to generalize effectively across the environment DAG.
A central obstacle in applying IRL to offline data, however, lies in the fact that the behavior policy underlying the dataset may not reflect expert behavior. To mitigate this, we introduce a rebalancing strategy inspired by [27] that biases trajectory sampling according to the reward of terminal states. Specifically, when estimating expectations under the expert policy $\mathbb{E}_{\pi_{E}}$, we sample trajectories $\tau=(s_{0},s_{1},\dots,s_{T})\in\mathcal{D}$ with probability $P(\tau)\propto R(s_{T})$, thereby constructing a rebalanced dataset that more closely approximates the visitation distribution of an expert policy. This reweighting adjusts edge visitation frequencies to favor high-reward regions, enabling more meaningful reward inference.
Inspired by the well-known maximum causal entropy IRL framework GAIL [28], we adopt the following adversarial training objective based on the rebalanced dataset:
@@
Nevertheless, the original GAIL framework induces a reward function that is strictly non-negative, making it inherently incapable of recovering the true underlying reward in our setting, which may include negative values (Equation 1). To address this limitation, we follow [29] and extract an unbiased edge-level reward function $R_{E}$ from the classifier:
@@
Intuitively, the learned edge reward reflects the degree to which an expert policy will prefer to traverse a given edge, thereby quantifying the edge’s importance for learning a high-quality GFlowNet policy. The complete edge reward extraction procedure is summarized in Algorithm 2 in Appendix A.
As shown in Figure 1(b) and Appendix D.2, the learned edge rewards exhibit strong generalization, assigning meaningful values even to transitions not observed in the dataset. This guides the policy to allocate attention across the entire environment DAG, rather than staying confined to regions near the training data, as in prior methods. Notably, such edge-level supervision aligns more closely with the original objective of GFlowNets: leveraging structured, multi-step decision-making to decompose the target distribution, rather than relying on proxy models for direct approximation. Next, we leverage these edge rewards to guide both DAG pruning and prioritized backward sampling of training trajectories, as illustrated in Figure 2.

### v2

# Beyond the Proxy: Trajectory-Distilled Guidance for Offline GFlowNet Training

@@ Abstract
Generative Flow Networks (GFlowNets) are effective at sampling diverse, high-reward objects, but in many real-world settings where new reward queries are infeasible, they must be trained from offline datasets. The prevailing proxy-based training methods are susceptible to error propagation, while existing proxy-free approaches often use coarse constraints that limit exploration. To address these issues, we propose Trajectory-Distilled GFlowNet (TD-GFN), a novel proxy-free training framework. TD-GFN learns dense, transition-level edge rewards from offline trajectories via inverse reinforcement learning to provide rich structural guidance for efficient exploration. Crucially, to ensure robustness, these rewards are used indirectly to guide the policy through DAG pruning and prioritized backward sampling of training trajectories. This ensures that final gradient updates depend only on ground-truth terminal rewards from the dataset, thereby preventing the error propagation. Experiments show that TD-GFN significantly outperforms a broad range of existing baselines in both convergence speed and final sample quality, establishing a more robust and efficient paradigm for offline GFlowNet training.
@@ 1 Introduction
Figure 1: Comparison of methods on the Molecule Design task using a $1,500$-trajectory dataset. Top-$10$ Reward is the average of the $10$ best samples, and Convergence is the number of training trajectories required for convergence. We compare against two proxy-based baselines from Bengio et al. (2021): Proxy-GFN (using a proxy learned on this dataset) and Oracle-GFN (using the oracle as a best-case proxy). See Section 4.3 for details.
@@
Generative Flow Networks (GFlowNets or GFNs) (Bengio et al., 2021; 2023) are a promising class of generative models designed to sample compositional objects in proportion to their positive rewards. They have demonstrated strong empirical success across diverse domains, including molecule discovery (Bengio et al., 2021), biological sequence design (Jain et al., 2022), combinatorial optimization (Zhang et al., 2025a), and the fine-tuning of generative models (Hu et al., 2024a; Liu et al., 2025). Despite their potential, in many real-world applications, interacting with the environment to obtain reward feedback is often impractical or infeasible due to costly experiments or the need for human evaluation. This necessitates learning GFlowNets from historically collected datasets.
The current prevailing approach involves training a proxy model on the dataset to approximate the reward signal, which the GFlowNet then queries during training to obtain rewards for the terminal states and update itself. However, constructing a high-quality reward model is resource-intensive, typically demanding a large and diverse dataset along with substantial domain-specific expertise (Ouyang et al., 2022; Jain et al., 2023a). Moreover, queries for samples that are out-of-distribution for the proxy, or those it estimates inaccurately, can propagate errors directly through gradients and corrupt the policy. This challenge is particularly pronounced in domains such as language modeling (Ziegler et al., 2019; Dalrymple et al., 2024), recommendation systems (Chen et al., 2024), and autonomous driving (Xie et al., 2024).
@@
Consequently, some recent efforts, such as RO-GFlowNets (Wang et al., 2023) and COFlowNet (Zhang et al., 2025b), have begun to explore abandoning the proxy to learn directly from offline trajectories. Such trajectory data commonly arise in domains where GFlowNets can be naturally applied, including molecular simulation (van der Kamp et al., 2010), robotics (Khazatsky et al., 2024), and symbolic reasoning (Cobbe et al., 2021). While often overlooked in proxy-based methods, this trajectory-level information is a cornerstone of offline reinforcement learning (Kumar et al., 2020; Kostrikov et al., 2021) that has yet to be fully exploited for GFlowNets. By learning directly from complete, pre-collected trajectories rather than relying on trial-and-error exploration, these methods train more efficiently than their proxy-based counterparts, as illustrated in Figure 1. However, they typically impose coarse constraints to align the policy with the dataset, a practice that can hinder generalization, lead to poor exploration, and is highly sensitive to data quality.
In contrast to imposing such naive constraints, we introduce Trajectory-Distilled GFlowNet (TD-GFN), a novel proxy-free framework that learns to extract rich structural information on the environment’s directed acyclic graph (DAG) to guide its policy. Our work is built on an insight that different edges in the DAG are not equally important for learning an effective GFlowNet policy (also noted in prior work, e.g., Silva et al. (2025), though in a different setting). For instance, as shown in Figure 2(a), removing the red edge $(n_{1}\rightarrow n_{2})$ blocks access to a terminal node with reward $20$, while removing the green edge $(n_{4}\rightarrow n_{7})$ only affects a low-reward node ($0.1$). To quantify this varying importance of edges, we employ inverse reinforcement learning (IRL) on a rebalanced offline dataset, leveraging the theoretical connection between GFlowNet training and reinforcement learning (Tiapkin et al., 2024) to distill transition-level scores we term edge rewards, as illustrated in Figure 2(b). These learned rewards provide dense, intermediate guidance that enables the policy to generalize across the DAG, fundamentally differing from a proxy model in both its learning objective and its method of application.
Crucially, to ensure robustness, we use these edge rewards in an effective indirect manner. First, we prune low-utility transitions as determined by their edge rewards, yielding a more compact and expressive action space that prioritizes high-value trajectories (Figure 2(c)). Second, we introduce a prioritized backward sampling procedure guided by both terminal and edge rewards, which strategically allocates the model’s attention across both terminal states and intermediate pathways during training. Throughout this process, the gradient-based updates to our policy rely only on the ground-truth terminal rewards from the dataset. This indirect use of the learned edge rewards insulates the policy from potential inaccuracies in the IRL phase, preventing the error propagation that plagues proxy-based methods. More importantly, the dense reward guidance throughout its decision-making process enables TD-GFN to explore high-reward samples with superior efficiency. As shown in Figure 1, it achieves performance competitive with, and even slightly surpassing the Oracle-GFN, a GFlowNet trained on a best-case proxy (pre-trained on $300k$ molecules) from Bengio et al. (2021), while using $20$ times fewer training trajectories.
Extensive experiments demonstrate that our proposed TD-GFN trains with superior efficiency and reliability, significantly outperforming a wide range of baselines in convergence speed and final sample quality. These results establish TD-GFN as a more robust, stable, and efficient paradigm for offline GFlowNet training, marking a significant step forward from conventional methods.
In summary, our key contributions are as follows:
  • A Novel Proxy-Free Paradigm: We introduce TD-GFN, a new paradigm for training GFlowNets from offline trajectory data that eliminates reliance on proxy models by avoiding out-of-distribution reward queries. It employs IRL to learn an edge reward for each transition on the DAG from a rebalanced offline dataset, which is then used to guide the policy’s training and generalization across the environment.
  • Robust Guidance Mechanism: We propose two mechanisms for indirectly guiding the policy using the learned edge rewards. Through pruning the DAG and prioritized backward sampling of training trajectories based on these edge rewards, we enable the model to train efficiently under dense guidance and ensure that the final gradient updates rely only on ground-truth rewards, thereby preventing the error propagation.
  • State-of-the-Art Performance and Efficiency: Through extensive experiments, we demonstrate that TD-GFN establishes a new state of the art for offline GFlowNet training. It significantly outperforms a wide range of baselines by robustly achieving a better fit to the target distribution, faster convergence, and higher sample quality.
@@ 2 Preliminaries
In this section, we describe Generative Flow Networks (GFlowNets or GFNs) and the maximum causal entropy inverse reinforcement learning (IRL) framework, which help clarify our proposed method. Additional related works are provided in Appendix B for completeness.
@@ 2.1 GFlowNets
GFlowNets (Bengio et al., 2021; 2023) are defined within an environment represented as a directed acyclic graph (DAG), denoted by $G=(V,E)$, where $V$ is the set of nodes and $E$ the set of directed edges. A node $s^{\prime}$ is said to be a child of node $s$ if and only if there exists a directed edge $(s\rightarrow s^{\prime})\in E$; correspondingly, $s$ is the parent of $s^{\prime}$. The graph includes a unique root node $s_{0}\in V$ with no incoming edges and a subset of terminal nodes $\mathcal{X}\subseteq V$ with no outgoing edges.
A positive reward function $R(x)$ is defined over the terminal nodes $x\in\mathcal{X}$. The objective of GFlowNets is to learn a stochastic forward policy $\mathcal{P}_{F}(s^{\prime}|s)$ such that the probability of reaching a terminal node $x\in\mathcal{X}$ is proportional to its associated reward $R(x)$. This is accomplished by generating sequences of transitions, where at each step, the agent selects a child node of the current state. In this work, we focus on deterministic environment dynamics, where each action uniquely determines the next state by traversing the corresponding edge in the DAG, as is standard in most existing GFlowNet studies (Jain et al., 2022; Jang et al., 2023; Hu et al., 2024a; Zhang et al., 2025a). Recent work (Tiapkin et al., 2024) establishes a formal equivalence between GFlowNet training and entropy-regularized RL under specific settings: a discount factor $\gamma=1$, an entropy regularization coefficient $\lambda=1$, and a modified reward function defined by a fixed backward policy $\mathcal{P}_{B}(s|s^{\prime})$ that samples parent nodes in the DAG in reverse from a given child node:
@@ 2.2 Maximum causal entropy IRL
Maximum causal entropy inverse reinforcement learning (Ziebart et al., 2008; 2010) seeks to recover a reward function $r(s,s^{\prime})$ that both explains expert behavior under a policy $\pi_{E}$ and promotes maximum causal entropy in the learned policy. This objective is formalized as:
$\underset{r}{\text{minimize}}\left(\max_{\pi}\left[\lambda\mathcal{H}(\pi)+\mathbb{E}_{\pi}\left(\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},s_{t+1})\right)\right]\right)-\mathbb{E}_{\pi_{E}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},s_{t+1})\right],$ (3)
where $\mathcal{H}(\pi)$ denotes the causal entropy of the policy $\pi$, consistent with its definition in Equation (2). In our approach, we adapt this framework to infer an edge-level reward function $R_{E}:E\rightarrow\mathbb{R}$ that reflects the expert policy’s preferences for allocating attention across specific edges in the graph.
@@ 3 Method
In this section, we introduce Trajectory-Distilled GFlowNet (TD-GFN), a novel proxy-free training framework. Unlike prior methods that constrain the policy to remain close to the dataset (Wang et al., 2023; Zhang et al., 2025b), TD-GFN achieves a more effective exploration–exploitation trade-off by strategically allocating attention across the environment DAG.
The overall training pipeline is illustrated in Figure 3. First, TD-GFN learns an edge reward function from a rebalanced dataset using inverse reinforcement learning (IRL). This function captures the relative importance of each edge in the environment DAG, providing principled guidance for policy learning. Leveraging these edge-level rewards, TD-GFN then prunes the DAG to remove low-utility transitions and trains the policy via prioritized backward sampling over the resulting subgraph. This design enables strategic generalization to broad, unseen regions of the state space while maintaining alignment with high-reward behaviors. The following subsections detail each component. A comprehensive summary of the full procedure is presented in Algorithm 1 (Appendix A).
@@ 3.1 Edge rewards extraction via IRL
Motivated by the insight that different edges in the environment DAG contribute unequally to effective policy learning, we aim to quantify this disparity to enable more targeted and generalizable training. Nevertheless, measuring the contribution of individual transitions under the inherently stochastic nature of GFlowNet policies poses a fundamental challenge. To tackle this, we leverage the theoretical equivalence between GFlowNet training and entropy-regularized reinforcement learning (Equation (2)) and applying the maximum causal entropy IRL framework (Ziebart et al., 2008; 2010) (Equation (3)) to extract a fine-grained edge-level reward function from offline trajectories. These learned edge rewards act as structured auxiliary signals that capture the relative importance of transitions, guiding the policy to generalize effectively across the environment DAG.
A central obstacle in applying IRL to offline data, however, lies in the fact that the behavior policy underlying the dataset may not reflect expert behavior. To mitigate this, we introduce a rebalancing strategy inspired by Hong et al. (2023) that biases trajectory sampling according to the reward of terminal states. Specifically, when estimating expectations under the expert policy $\mathbb{E}_{\pi_{E}}$, we sample trajectories $\tau=(s_{0},s_{1},\dots,s_{T})\in\mathcal{D}$ with probability $P(\tau)\propto R(s_{T})$, thereby constructing a rebalanced dataset that more closely approximates the visitation distribution of an expert policy. This reweighting scheme effectively creates a pseudo-expert dataset by adjusting edge visitation frequencies to favor high-reward regions, thereby aligning the learned edge rewards with expert behavior more closely.
Inspired by the well-known maximum causal entropy IRL framework GAIL (Ho & Ermon, 2016) which has been theoretically and empirically shown to effectively learn reward functions with strong generalization (Xu et al., 2020; Luo et al., 2022; 2024), we adopt the following adversarial training objective based on the rebalanced dataset:
@@
Nevertheless, the original GAIL framework induces a reward function that is strictly non-negative, making it inherently incapable of recovering the true underlying reward in our setting, which may include negative values (Equation (1)). To address this limitation, we follow Kostrikov et al. (2018) and extract an unbiased edge-level reward function $R_{E}$ from the classifier:
@@
Intuitively, the learned edge reward reflects the preference of the near-expert behavior policy underlying the rebalanced dataset, thereby quantifying the edge’s importance for learning a high-quality GFlowNet policy. Fundamentally, this differs from proxy reward models: it is not designed to predict the reward of each terminal state, but rather to provide indirect, non-gradient guidance to the policy, as we demonstrate in the following section. The complete edge reward extraction procedure is summarized in Algorithm 2 in Appendix A.
As shown in Figure 2(b) and Appendix D.3, the learned edge rewards exhibit strong generalization, assigning meaningful values even to transitions not observed in the dataset. This guides the policy to allocate attention across the entire environment DAG, rather than staying confined to regions near the training data, as in prior methods. Notably, such edge-level guidance across the entire DAG aligns more closely with the GFlowNet philosophy: leveraging structured, multi-step decision-making to decompose the complex problem of sampling from a target distribution, rather than attempting to model the distribution directly.